OpenStudy (anonymous):

A group of backpackers started with 5 pounds of cheese. On the second day they had only 2 1/2 pounds. On the third day they had only 1 1/4pounds. If the pattern continues, how much will they have on the fourth day?

OpenStudy (anonymous):

they are eating half of the chees each day 5 , 5*1/2 . 5*1/2* 1/2 so what do you think they have on 4 th day?

OpenStudy (anonymous):

do you see the pattern? day 1 5 day 2 2.5 day 3 1.25 day 4 1.25 / 2

OpenStudy (anonymous):

its 1.25 pounds of cheese divdied by 2

OpenStudy (anonymous):

or if you want it in fractions its 1 1/4 / 2

OpenStudy (anonymous):

yes i see it . thank you so much.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

as a fraction we get 1 /1/4 = 5/4 5/4 / 2 = 5/4 * 1/2 = 5/8 pounds cheese